OpenStudy (anonymous):

Solve for x. (3/x)+(2/a)=(1/b)

jhonyy9 (jhonyy9):

3/x + 2/a =1/b 3ab+2bx=ax x(2b-a)=-3ab x=-3ab/(2b-a)